### v1

@@ III-C2 Smart Walking Aid - Haptic Feedback and Motion Data
The ESP32 MCU receives processed gait event data from the SFO via wireless communication using either BLE or Wi-Fi. The firmware was developed in C/C++ using the Arduino integrated development environment (IDE) and enables the control of two actuators to enable vibration patterns depending on pre-defined overloads. Based on predefined relative thresholds, derived from the user’s body weight, the received plantar pressure values from the SFO are used to trigger both vibration actuators. For instance, if the plantar pressure during the stance phase exceeds 120% of the expected load distribution, a feedback alert can be activated. However, such thresholds are individual to each patient and should be defined by a medical expert [105]. Vibration signals are emitted using a configurable pulse-based pattern. Depending on the detected parameter both the intensity and interval of the vibration can be dynamically adapted. This customization allows targeted feedback for specific gait abnormalities, such as in-toeing or out-toeing patterns, where intensity can be increases with foot rotation angles. The integrated IMU, initialized with the same configuration as in the SFO, continuously records motion data for activity recognition. This data is also streamed to the smartphone application. To extend the functionality, these data could be used in future implementations for activity classification or fall detection.
III-C3 Smartphone Application
@@
The design of the smartphone application builds upon the main functions proposed by Resch et al. [64] for a smart insole app dedicated to foot health monitoring. Building on these proposed functions, our application implements key features such as a heatmap visualization of plantar pressure based on the FSR sensors, real-time gait characteristics, and a walking diary displaying step counts and activity-related health status ratings. Additional modules include a knowledge database with a search function for articles on foot health, a training area featuring video tutorials for foot-related exercises, and a map for locating local medical specialists. The app also enables monitoring of the communication status between the smart insole and the walking aid.
For pressure data visualization, a custom heatmap fragment shader was implemented. Its output was based on the number of FSRs used and colored using the Turbo colormap developed by Google  [^4] https://research.google/blog/turbo-an-improved-rainbow-colormap-for-visualization/. The application was developed in Android Studio  [^5] https://developer.android.com/studio (version 2024.2.1) using Kotlin as the main programming language. We utilized Jetpack Compose  [^6] https://developer.android.com/compose and its integrated UI libraries from Material2 and Material3, enabling flexible design of forms, buttons, and icons. Wireless communication was implemented via Wi-Fi, Bluetooth, and MQTT protocols to ensure real-time data transmission between devices. The prototype was deployed as an android package kit (APK) on an Android 14 device using software development kit (SDK) version 35. Figure 5 shows selected screenshots of the main functionalities of the developed mHealth application.
III-D Prototype Validation Study: Experimental Setup
We conducted an experimental user study with eight participants to evaluate the integrated system using both objective and subjective quantitative measurements. The study comprised three main validation objectives:
@@ III-D Prototype Validation Study: Experimental Setup
    Data Accuracy: Validation of the SFOs plantar pressure and spatial/temporal gait data against an instrumented treadmill.
@@
    System Acceptance and Feedback Perception: Evaluation of two feedback mechanisms using a customized questionnaire.
@@
    Usability Evaluation: User experience evaluation of the smartphone app based on the System Usability Scale (SUS[106].
In the following subsections, the study design, apparatus, procedure and tasks, as well as the applied measures and data analysis are documented.
III-D1 Study Design
@@
Figure 6: Visual overview of the experimental validation process including orthosis testing, haptic feedback evaluation, and usability assessment.
@@
To evaluate the influence of haptic feedback on user perception, participants tested both the SFO and the crutch, each equipped with one vibration module. The study was based on a $2\times 2$ factorial within-subject design with two independent variables: System (two levels: Orthosis and Crutch) and Feedback Type (two levels: Continuous and Pulsed Pattern). Dependent variables were based on subjective ratings regarding noticeability, intrusiveness, regular usage, and perceived usefulness. A $2\times 2$ balanced Latin square design was applied to counterbalance the conditions and avoid sequence effects [107].
Finally, the smartphone app was tested to assess perceived usability using the standardized SUS questionnaire.
III-D2 Apparatus
@@
For the evaluation of motion and pressure data acquired from the IMU and FSR sensors, we used an instrumented treadmill, which is considered the gold standard for validating gait analysis systems. Specifically, a RehaWalk® FDM-THPL-S-2i instrumented treadmill (Zebris Medical GmbH) was used to measure spatial and temporal gait parameters. The system is based on an h/p/cosmos pluto treadmill and integrates 3,120 capacitive sensors ($\pm$ 0.05% measurement accuracy), arranged in an 80 $\times$ 39 grid layout, covering a sensor area of 101.6 × 49.5 cm. Data were recorded at a sampling rate of 100 Hz (similar to the SFO), within a sensor range of 1–120 N/cm². Zebris FDM software (version 2.0.4) was used to run gait analysis and export CSV raw data. The software ran on a Windows 10 Pro system with an AMD Ryzen 5900X 12-core processor 3.70 GHz, a GeForce RTX 3070 graphics card, and 32 GB RAM.
III-D3 Procedure and Tasks
@@
Prior to study participation, all individuals completed a brief pre-screening to confirm that these criteria were met. The study was conducted under controlled laboratory conditions without external influences.
First, participants were informed about the study procedures and that they could withdraw or discontinue participation at any time. After providing informed consent, they completed a survey about their demographic data, foot conditions, previous experience with assistive devices, and mHealth applications. Furthermore, participants were asked to rate their familiarity with the use of smartphone healthcare apps.
Following this, the SFO was provided to each participant as a fully preassembled system. An integrated vacuum-based mechanism was used to ensure optimal adaptation to the individual leg shape. After a brief instruction, each participant completed five walking trials on the Zebris treadmill at a fixed speed of 1.5 km/h. Each trial lasted 10 seconds to ensure the recording of at least one full gait cycle, typically capturing multiple cycles for averaging and comparison across multiple measurements. Participants were instructed to walk as normally as possible without holding onto the treadmill handles.
After the walking trials, the walking aid prototype was introduced and tested both while standing and walking, in combination with the SFO. Haptic feedback was evaluated for both systems under predefined overload conditions to demonstrate feedback alerts. Two feedback type modalities (continuous and pulsed patterns) were activated in randomized order for each participant. Following this, participants completed a questionnaire consisting of quantitative Likert-items to assess noticeability, intrusiveness, regular usage, and perceived usefulness. The questionnaire items were as follows:
  • Noticeability: ”The haptic feedback was clearly noticeable during use.
  • Non-Intrusiveness ”I found the feedback comfortable and non-intrusive.
  • Regular Usage ”I could imagine using such feedback during everyday mobility tasks.
  • Usefulness ”The feedback would be useful in real-life situations (e.g., alerting me of incorrect foot placement or walking posture).
Additionally, participants were asked to indicate their preferred vibration alert type in a single-choice question.
III-D4 Measurement and Data Analysis
A quantitative measurement approach was chosen to evaluate the overall system and its individual subsystems. Descriptive and inferential statistical analyses were conducted using R (version 4.2.2) with the rstatix package [108]. The following paragraphs describe the corresponding measurements and data analysis in detail.
Smart Orthosis: Validation of Plantar Pressure and Gait Parameters
Both systems recorded gait events and pressure-related data simultaneously. Manual triggering was performed during the swing phase of the right leg (foot orthosis) to synchronize both systems at the moment of initial right foot contact. Each trial lasted for 10 seconds, recorded at a sampling rate of 100 Hz in both systems, resulting in five datasets per participant. The Zebris FDM software enabled the export of pressure data and gait characteristics were aggregated as mean and standard deviation values per trial. The definitions of the respective gait parameters are provided in Table I.
TABLE I: Overview of analyzed gait parameters grouped by parameter type.
Parameter Type Parameter Unit
Temporal Stance Duration s
Swing Duration s
Cycle Duration s
Spatial Step Length m
Cycle Length m
Kinetic Mean Plantar Pressure N/cm2
For pressure data the full time-series dataset was used for the SFO, while the treadmill provided aggregated data for one step. However, a direct quantitative comparison of force or pressure values between the treadmill and orthosis systems is not possible, due to fundamental differences in sensor configuration and measurement principles. The treadmill measures plantar pressure at the interface between the orthosis sole and the treadmill surface, while the orthosis records forces between the foot and the insole. As a result, the absolute values and spatial resolution of the pressure data differ between systems. Therefore, pressure data were visualized as heatmaps and force–time curves to qualitatively assess whether gait events could be reliably identified from the force signal progression.
Smart Orthosis and Walking Aid: Haptic Feedback Assessment
To assess the user experience and perceived effectiveness of the haptic feedback system, the participants rated four statements on a 7-point Likert scale [109]. The items addressed noticeability, intrusiveness, regular usage, and perceived usefulness, with response options ranging from 1 (strongly disagree) to 7 (strongly agree). Descriptive statistics were applied to summarize the participants’ responses and identify trends in agreement levels. A two-factorial repeated-measures (RM) Analysis of Variance (ANOVA) was used to assess the effect of system and feedback type on noticeability, intrusiveness, regular usage, and perceived usefulness.
Smartphone Application: Usability Evaluation
Usability of the smartphone application was assessed using the standardized SUS, which is frequently used to assess user experience of interactive systems such as smartphone apps. The questionnaire consists of ten items rated on a five-point Likert scale ranging from 1 (strongly disagree) to 5 (strongly agree). Even-numbered items were reverse-scored and calculated according to the specifications by Brooke [106]. The total score was calculated and converted into a value between 0 and 100 as a quantifiable measure of overall usability. Descriptive statistics were used to analyze the SUS scores.
IV Results
The results section begins with an overview of the study group characteristics (Section IV-A). This is followed by a comparative evaluation of the accuracy data obtained from both the SFO and the instrumented treadmill (Section IV-B). Subsequently, the evaluation of the haptic feedback functionality of the forearm crutch and the SFO is presented (Section IV-C). Finally, the usability evaluation results of the accompanying smartphone application are presented (Section IV-D).
IV-A Participants
@@
In total, 8 participants (4 female, 4 male) were recruited via mailing lists from our institution. The age of the participants ranged between 24 and 72 years (M = 39.75, SD = 17.92). Participants’ educational backgrounds included four with vocational training, three with a bachelor’s degree, and one with a master’s degree. Occupations were distributed as follows: services and sales (n = 2), engineering (n = 2), other fields (n = 2), clerical (n = 1), and health-related professions (n = 1). The shoe size ranged from EU 38 to EU 43 (M = 40.37, SD = 1.85). All participants reported having a current or past foot condition. Four had previous conditions (such as a broken leg, sprained ankle, pes planus, or torn ligament). Two participants reported chronic conditions (foot malalignment and arthrosis), one participant had an acute condition (sprained ankle), and one reported both an acute (sprained ankle) and chronic condition (torn ligament and previous leg fracture). All participants indicated that they had previously used orthopedic devices, including insoles (n = 7), ankle braces or bandages (n = 3), and foot orthoses (n = 2). Furthermore, four participants already used forearm crutches as medical walking aids. In total, three participants have previous experience with a mobile health app. Participants rated their familiarity with the use of smartphone healthcare apps on a scale from 1 (not at all familiar) to 5 (extremely familiar), with an average rating of 2.38 (SD = 1.60). All participants completed the study and were included in the analysis.
@@
IV-B Smart Orthosis: Objective Accuracy of Plantar Pressure and Gait Parameters
IV-B1 Gait Data Measurement
@@
To assess the reliability of temporal and spatial gait parameters, the accuracy of the SFO measurements was evaluated against the treadmill system across all trials. The results showed that temporal parameters achieved mean accuracies above 85%, except for swing phase duration, which reached only 60%. For spatial parameters, mean accuracies were at least 81%, although with higher standard deviations. The descriptive results are summarized in Table II.
@@
TABLE II: Mean and standard deviation of gait parameter accuracy of the SFO compared to treadmill measurements across all participants.
Parameter Type Mean Accuracy SD Accuracy
Temporal
Stance Phase Duration 96.7% $\pm$ 3.1
@@
Spatial
Step Length 81.0% $\pm$ 12.9
Cycle Length 83.4% $\pm$ 13.3
Figure 7: Top: Measured plantar pressure data of Trial 4 (P3), showing force values recorded by multiple FSR sensors over the duration of the task. Distinct gait phases can be identified from the characteristic patterns in the force signals. The left axis represents the force values per sensor, while the right axis indicates the total force (black line) across all included sensors. Bottom: Measured acceleration and gyroscope data from the IMU for each of the three axes, illustrating gait phase conformity with the recorded pressure data.
@@
The Shapiro–Wilk normality test indicated that both cycle phase duration ($p>.095$) and swing phase duration ($p>.501$) were normally distributed, while stance phase duration was not ($p=.013$).
A paired t-test revealed a statistically significant difference between SFO and treadmill for swing phase duration ($t(7)=9.74$, $p<.001$), Cohen’s d = 3.44 (large effect size), as well as for cycle phase duration (t(7)=18.7t(7)=-18.7, $p<.001$), Cohen’s d = 6.59 (large effect size), which was significantly shorter in SFO compared to treadmill. Furthermore, a Wilcoxon signed-rank test on stance duration data showed no significant difference ($W=8$, $p=.195$, with moderate effect $r=0.50$).
A Pearson correlation analysis showed a strong and significant correlation for cycle duration ($r=0.973$, $p<.001$), and a moderate but not statistically significant correlation for swing duration ($r=0.673$, $p=.067$). Additionally, a Spearman correlation analysis for stance duration, which was not normally distributed, did not reveal a significant correlation ($\rho=-0.167$, $p=.693$).
@@
The Shapiro–Wilk test indicated that both step length ($p>.091$) and cycle length ($p=.259$) were normally distributed.
A paired t-test revealed a statistically significant difference in step length between the treadmill and orthosis conditions (t(7)=4.38t(7)=-4.38, $p=.003$), Cohen’s d = 1.55 (large effect size), indicating that step length was significantly shorter when measured with the orthosis compared to the treadmill. Furthermore, a statistically significant difference was found for cycle length (t(7)=3.65t(7)=-3.65, $p=.008$), Cohen’s d = 1.29 (large effect), with the orthosis also leading to lower values.
Pearson correlation analysis showed a statistically significant large correlation for step length ($r=0.760$, $p=.029$). However, no significant correlation was found for cycle length ($r=0.678$, $p=.065$).
IV-B2 Pressure Data Measurement

### v2

@@ 3.3.2 Smart Walking Aid—Haptic Feedback and Motion Data
The ESP32 MCU receives processed gait event data from the SFO via wireless communication using either BLE or Wi-Fi. The firmware was developed in C/C++ using the Arduino integrated development environment (IDE) and enables the control of two actuators to generate vibration patterns depending on pre-defined overloads. In the current prototype implementation, relative pressure thresholds were defined as fixed percentage values of the user’s body weight to ensure consistent triggering conditions during system testing, similar to Mariani et al. Mariani et al. (2013), who used a threshold of 5% of body weight for gait phase detection. This constant threshold design served to demonstrate the feedback functionality under controlled conditions and was not intended to represent clinically calibrated loads. In future implementations, these thresholds could be dynamically personalized and adapted per gait phase (stance vs. swing) based on baseline load measurements obtained during standing and normal walking. However, such thresholds are individual to each patient and thresholds should be defined and validated by medical experts to ensure the detection of clinically relevant overload conditions Sardini et al. (2015). Vibration signals are emitted using a configurable pulse-based pattern. Depending on the detected parameter both the intensity and interval of the vibration can be dynamically adapted. This customization allows targeted feedback for specific gait abnormalities, such as in-toeing or out-toeing patterns, where intensity can be increased with foot rotation angles. The integrated IMU, initialized with the same configuration as in the SFO, continuously records motion data for activity recognition. This data is also streamed to the smartphone application. To extend the functionality, these data could be used in future implementations for activity classification or fall detection.
3.3.3 Smartphone Application
The design of the smartphone application builds upon the main functions proposed by Resch et al. Resch et al. (2023) for a smart insole app dedicated to foot health monitoring. Building on these proposed functions, our application implements key features such as a heatmap visualization of plantar pressure based on the FSR sensors, real-time gait characteristics, and a walking diary displaying step counts and activity-related health status ratings. Additional modules include a knowledge database with a search function for articles on foot health, a training area featuring video tutorials for foot-related exercises, and a map for locating local medical specialists. The app also enables monitoring of the communication status between the SFO and the walking aid.
For pressure data visualization, a custom heatmap fragment shader was implemented. Its output was based on the number of FSRs used and colored using the Turbo colormap developed by Google (2019) (Google). The application was developed in Android Studio Studio (version 2024.2.1) using Kotlin as the main programming language. We utilized Jetpack Compose Developers and its integrated UI libraries from Material2 and Material3, enabling flexible design of forms, buttons, and icons. Wireless communication was implemented via Wi-Fi, Bluetooth, and MQTT protocols to ensure real-time data transmission between devices. The prototype was deployed as an android package kit (APK) on an Android 14 device using software development kit (SDK) version 35. Figure 5 shows selected screenshots of the main functionalities of the developed mHealth application.
@@ 3.3.3 Smartphone Application
3.4 Prototype Validation Study: Experimental Setup
A technical feasibility and validation study was conducted with eight participants to evaluate the integrated system in terms of both objective measurement accuracy and subjective user experience. The primary goal of this proof-of-concept investigation was to verify the system’s functional reliability and multimodal feedback performance under controlled laboratory conditions, rather than to focus on algorithmic novelty or benchmark testing of individual algorithms, or to conduct a clinical study. The study comprised two supporting validation objectives (system feasibility and application usability) and one primary objective (feedback evaluation).
@@ 3.4 Prototype Validation Study: Experimental Setup
    System Feasibility: Validation of the SFO plantar pressure and spatial/temporal gait data against an instrumented treadmill.
@@
    Application Usability: User experience evaluation of the smartphone app based on the System Usability Scale (SUSBrooke et al. (1996).
@@
    Feedback Evaluation: Assessment of two feedback mechanisms applied within the SFO and forearm crutch using a customized questionnaire.
In the following subsections, the study design, apparatus, procedure and tasks, as well as the applied measures and data analysis, are documented.
3.4.1 Study Design
@@
Figure 6: Visual overview of the experimental validation process including orthosis testing, usability assessment, and haptic feedback evaluation.
@@
The smartphone app was tested to assess perceived usability using the standardized SUS questionnaire.
To evaluate the influence of haptic feedback on user perception, participants tested both the SFO and the crutch, each equipped with one vibration module. The study was based on a $2\times 2$ factorial within-subject design with two independent variables: System (two levels: Orthosis and Crutch) and Feedback Type (two levels: Continuous and Pulsed Pattern). Dependent variables were based on subjective ratings regarding noticeability, intrusiveness, regular usage, and perceived usefulness. A $2\times 2$ balanced Latin square design was applied to counterbalance the conditions and avoid sequence effects Schwind et al. (2023).
3.4.2 Apparatus
@@
For the evaluation of motion and pressure data acquired from the IMU and FSR sensors, we used an instrumented treadmill, which is considered the gold standard for validating gait analysis systems. Specifically, a RehaWalk® FDM-THPL-S-2i instrumented treadmill (Zebris Medical GmbH, Isny, Germany) was used to measure spatial and temporal gait parameters. The system is based on an h/p/cosmos pluto treadmill and integrates 3120 capacitive sensors ($\pm$0.05% measurement accuracy), arranged in an 80 $\times$ 39 grid layout, covering a sensor area of 101.6 × 49.5 cm. Data were recorded at a sampling rate of 100 Hz (similar to the SFO), within a sensor range of 1–120 N/cm². Zebris FDM software (version 2.0.4) was used to run gait analysis and export CSV raw data. The software ran on a Windows 10 Pro system with an AMD Ryzen 5900X 12-core processor, 3.70 GHz, a GeForce RTX 3070 graphics card, and 32 GB RAM.
3.4.3 Procedure and Tasks
The study was conducted under controlled laboratory conditions without external influences. First, participants were informed about the study procedures and that they could withdraw or discontinue participation at any time. After providing informed consent, they completed a survey about their demographic data, foot conditions, previous experience with assistive devices, and mHealth applications. Furthermore, participants were asked to rate their familiarity with the use of smartphone healthcare apps. Following this, the SFO was provided to each participant as a fully preassembled system. An integrated vacuum-based mechanism was used to ensure optimal adaptation to the individual leg shape. After a brief instruction, each participant completed five walking trials on the Zebris treadmill at a fixed speed of 1.5 km/h. Each trial lasted 10 s to ensure the recording of at least one full gait cycle, typically capturing multiple cycles for averaging and comparison across multiple measurements. Participants were instructed to walk as normally as possible without holding onto the treadmill handles.
After the walking trials, the walking aid prototype was introduced and tested both while standing and walking, in combination with the SFO. Haptic feedback was evaluated for both systems under predefined overload conditions to demonstrate feedback alerts and to ensure identical vibration triggering conditions across all participants. Two feedback type modalities (continuous and pulsed patterns) were activated in randomized order for each participant. Following this, participants completed a questionnaire consisting of quantitative Likert items to assess noticeability, intrusiveness, regular usage, and perceived usefulness. Additionally, participants were asked to indicate their preferred vibration alert type in a single-choice question.
3.4.4 Measurement and Data Analysis
A quantitative measurement approach was chosen to evaluate the overall system and its individual subsystems. Descriptive and inferential statistical analyses were conducted using R (version 4.2.2) with the rstatix package Kassambara (2023). The following paragraphs describe the corresponding measurements and data analysis in detail.
System Feasibility: Accuracy of Plantar Pressure and Gait Parameters (Smart Orthosis)
Both systems recorded gait events and pressure-related data simultaneously. Manual triggering was performed during the swing phase of the right leg (foot orthosis) to synchronize both systems at the moment of initial right foot contact. Each trial lasted for 10 s, recorded at a sampling rate of 100 Hz in both systems, resulting in five datasets per participant. The Zebris FDM software enabled the export of pressure data and gait characteristics were aggregated as mean and standard deviation values per trial. The definitions of the respective gait parameters are provided in Table 1.
Table 1: Overview of analyzed gait parameters grouped by parameter type.
Parameter Type Parameter Unit
Temporal Stance Duration s
Swing Duration s
Cycle Duration s
Spatial Stride/Cycle Length m
Kinetic Mean Plantar Pressure N/cm2
For pressure data, the full time-series dataset was used for the SFO, while the treadmill provided aggregated data for one step. However, a direct quantitative comparison of force or pressure values between the treadmill and orthosis systems is not possible, due to fundamental differences in sensor configuration and measurement principles. The treadmill measures plantar pressure at the interface between the orthosis sole and the treadmill surface, while the orthosis records forces between the foot and the insole. As a result, the absolute values and spatial resolution of the pressure data differ between systems. Therefore, pressure data were visualized as heatmaps and force–time curves to qualitatively assess whether gait events could be reliably identified from the force signal progression.
Application Usability: Evaluation of Smartphone App Usability
Usability of the smartphone application was assessed using the standardized SUS, which is frequently used to assess user experience of interactive systems such as smartphone apps. The questionnaire consists of ten items rated on a five-point Likert scale ranging from 1 (strongly disagree) to 5 (strongly agree). Even-numbered items were reverse-scored and calculated according to the specifications by Brooke Brooke et al. (1996). The total score was calculated and converted into a value between 0 and 100 as a quantifiable measure of overall usability. Descriptive statistics were used to analyze the SUS scores.
Feedback Evaluation: User Perception of Haptic Feedback (Orthosis and Walking Aid)
To assess the user experience and perceived effectiveness of the haptic feedback system, the participants rated four statements on a 7-point Likert scale Likert (1932). The items addressed noticeability, intrusiveness, regular usage, and perceived usefulness, with response options ranging from 1 (strongly disagree) to 7 (strongly agree).
The questionnaire items included:
  • Noticeability: “The haptic feedback was clearly noticeable during use.
  • Non-Intrusiveness “I found the feedback comfortable and non-intrusive.
  • Regular Usage “I could imagine using such feedback during everyday mobility tasks.
  • Usefulness “The feedback would be useful in real-life situations (e.g., alerting me of incorrect foot placement or walking posture).
Descriptive statistics were applied to summarize the participants’ responses and identify trends in agreement levels. A two-factorial repeated-measures (RM) Analysis of Variance (ANOVA) was used to assess the effect of system and feedback type on noticeability, intrusiveness, regular usage, and perceived usefulness.
3.5 Participants
@@
Prior to study participation, all individuals completed a brief pre-screening to confirm that these criteria were met. In total, 8 participants (4 female, 4 male) were recruited. The age of the participants ranged between 24 and 72 years (M = 39.75, SD = 17.92). Participants’ educational backgrounds included four with vocational training, three with a bachelor’s degree, and one with a master’s degree. Occupations were distributed as follows: services and sales (n = 2), engineering (n = 2), other fields (n = 2), clerical (n = 1), and health-related professions (n = 1). The shoe size ranged from EU 38 to EU 43 (M = 40.37, SD = 1.85). All participants reported having a current or past foot condition. Four had previous conditions (such as a broken leg, sprained ankle, pes planus, or torn ligament). Two participants reported chronic conditions (foot malalignment and arthrosis), one participant had an acute condition (sprained ankle), and one reported both an acute (sprained ankle) and chronic condition (torn ligament and previous leg fracture). All participants indicated that they had previously used orthopedic devices, including insoles (n = 7), ankle braces or bandages (n = 3), and foot orthoses (n = 2). Furthermore, four participants already used forearm crutches as medical walking aids. In total, three participants had previous experience with a mobile health app. Participants rated their familiarity with the use of smartphone healthcare apps on a scale from 1 (not at all familiar) to 5 (extremely familiar), with an average rating of 2.38 (SD = 1.60). All participants completed the study and were included in the analysis.
4 Results
The results section first presents the comparative accuracy validation between the SFO and the instrumented treadmill (Section 4.1), followed by the usability assessment of the smartphone application (Section 4.2), and concludes with the evaluation of the haptic feedback of both the orthosis and the walking aid (Section 4.3).
@@ 4 Results
4.1 System Feasibility: Accuracy of Plantar Pressure and Gait Parameters (Smart Orthosis)
4.1.1 Gait Data Measurement
@@
To assess the reliability of temporal and spatial gait parameters, the accuracy of the SFO measurements was evaluated against the treadmill system across all trials. The results showed that temporal parameters achieved mean accuracies above 85%, except for swing phase duration, which reached only 60%. For the spatial parameter, the mean accuracy was 83.4%, despite a high standard deviation. The descriptive results are summarized in Table 2.
@@
Table 2: Mean and standard deviation of gait parameter accuracy of the smart foot orthosis (SFO) compared to treadmill measurements across all participants.
Parameter Type Parameter Mean Accuracy SD Accuracy
Temporal Stance Phase Duration 96.7% $\pm$3.1
@@
Spatial Stride/Cycle Length 83.4% $\pm$13.3
@@
The Shapiro–Wilk normality test indicated that both cycle phase duration ($p>0.095$) and swing phase duration ($p>0.501$) were normally distributed, while stance phase duration was not ($p=0.013$). A paired t-test revealed a statistically significant difference between SFO and treadmill for swing phase duration ($t(7)=9.74$, $p<0.001$), Cohen’s d = 3.44 (large effect size), as well as for cycle phase duration (t(7)=18.7t(7)=-18.7, $p<0.001$), Cohen’s d = 6.59 (large effect size), which was significantly shorter in SFO compared to treadmill. Furthermore, a Wilcoxon signed-rank test on stance duration data showed no significant difference ($W=8$, $p=0.195$, with moderate effect $r=0.50$).
A Pearson correlation analysis showed a strong and significant correlation for cycle duration ($r=0.973$, $p<0.001$), and a moderate but not statistically significant correlation for swing duration ($r=0.673$, $p=0.067$). Additionally, a Spearman correlation analysis for stance duration, which was not normally distributed, did not reveal a significant correlation ($\rho=-0.167$, $p=0.693$).
@@
The Shapiro–Wilk test indicated that cycle length ($p=0.259$) was normally distributed. A paired t-test revealed a statistically significant difference in cycle length between the treadmill and orthosis conditions (t(7)=3.65t(7)=-3.65, $p=0.008$), Cohen’s $d=1.29$ (large effect), indicating that cycle length was significantly shorter when measured with the orthosis compared to the treadmill. Pearson correlation analysis showed no significant correlation for cycle length ($r=0.678$, $p=0.065$).
4.1.2 Pressure Data Measurement
Nevertheless, to compare both systems a qualitative comparison of plantar pressure distributions and force curves was conducted using visualizations from both systems for a representative trial (see Figure 8). The Zebris data represent averaged force curves for the three foot zones summarized from multiple gait cycles, whereas the orthosis data focus on one representative gait cycle from the same trial, which is also displayed in Figure 7.
Figure 7: (Top): Measured plantar pressure data of Trial 4 (P3), showing force values recorded by multiple FSR sensors over the duration of the task. Distinct gait phases can be identified from the characteristic patterns in the force signals. The left axis represents the force values per sensor, while the right axis indicates the total force (black line) across all included sensors. (Bottom): Measured acceleration and gyroscope data from the IMU for each of the three axes, illustrating gait phase conformity with the recorded pressure data.
Although absolute force values are not directly comparable due to differences in sensor configuration, both systems show similar temporal trends across the three anatomical foot zones and in the overall force–time diagram. Compared to the Zebris treadmill, which provides only aggregated contact data, the SFO offers step-wise heatmap visualizations, enabling a detailed analysis of the internal plantar pressure distribution for each gait event.